# UDEEP: Edge-based Computer Vision for In-Situ Underwater Crayfish and Plastic Detection

### v1

@@ 3 Methodology
The Methodology section is divided into four parts. The first part introduces the UDEEP platform workflow and followed by a review of the hardware setup. The third section explains the YOLO v5n, YOLO v5s, YOLO v8n and YOLO v8s model architectures and the final section reviews the dataset and delves into the data augmentation techniques applied to the crayfish and underwater plastic datasets.
3.1 UDEEP Platform Workflow
The UDEEP platform workflow involves two distinct steps. Initially, we categorise the dataset into different sets based on the objects of interest, crayfish or plastic. Subsequently, we train the system using CNN with these grouped images, encompassing various objects of interest. In the second step, during testing, we employ the Gaussian Mixture Model (GMM) to segment the testing images. Each segment was tested to determine if it corresponds to any of the specified objects of interest. Figure 1 below illustrates the proposed UDEEP workflow.
Figure 1: The proposed UDEEP platform workflow.
3.1.1 Training
The training phase is the initial step of any classification algorithm. This is undertaken on a given set of images that have been selected due to objects of interest being present in them. For the UDEEP platform, the images contained crayfish and plastic debris in the underwater environment. This phase is further divided into two sub-phases namely Segregation and feature generation.
3.1.2 Segregation
The images used in this Research project were sourced from rivers within the United Kingdom and contain crayfish and contain crayfish and plastics. CNNs don’t perform well in extracting features from complex scenes, they however excel in learning features when training images showcase single objects with minimal noise. However, in multi-object images, CNNs attempt to learn features collectively, making it challenging to precisely describe individual objects. The bounding box technique was employed to delineate and select the object(s) of interest in each image via the Roboflow platform Roboflow (2023). A bounding box serves as the minimum enclosing box around a specific object, capturing the entire object of interest while minimising noise. All training images are resized to a $416\times 416$ pixels.
3.1.3 Testing
To assess an underwater image, it’s essential to identify the various objects within it. The authors employed the GMM segmentation to determine the number of objects in an image, the classifications, and their outline. Subsequently, these segmented objects are fed into the trained CNN model to validate whether a given object is of interest or not.
3.1.4 segmentation
The task of image segmentation can be expressed as categorising pixels into semantic labels (semantic segmentation), delineating individual objects (instance segmentation), or a combination of both (panoptic segmentation). Semantic segmentation involves labelling pixels at the pixel level with specific object categories (such as crayfish or plastic) for all pixels in the image. Minaee et al. (2021)
GMM-based models, which consider pixel intensities, are employed for improving the robustness to Gaussian noise. The GMM method categorises an input image matrix into segments using the complete data log-likelihood concept. Each pixel’s Gaussian density function, multiplied by the probability of belonging to a segment, yields the log-likelihood based on GMM-generated segments. Similar to other segmentation methods, GMM faces the challenge of selecting the number of segments, commonly addressed by analysing the image histogram.
3.2 UDEEP underwater monitoring prototype: AI inference at the edge
@@
A meticulously designed and implemented prototype for underwater monitoring is illustrated in Figure 2. The system features an NJN equipped with an Intel RealSense D435i depth sensor. The NJN is used for accelerating the YOLO v5 and YOLO v8 inference at the edge. The NJN selected is tailored for diverse tasks, including image identification, object detection, data segmentation, and speech processing. It encompasses a robust configuration, boasting a 128-core Maxwell GPU, a quad-core ARM A57 Application Processor Unit (APU) running at 1.43 GHz, and 4 GB of 64-bit LPDDR4 memory operating at a speed of 25.6 GB/s. This system is further enhanced by the inclusion of an NJN, preloaded with a pre-trained DL model.
Figure 2: UDEEP monitoring system.
@@ 3.3 Network Architecture
Four object classification network architectures were used from the YOLO family used in this research project, driven by the necessity to perform object detection at the edge where processing and power resources are constrained. The choice leaned towards the compact and efficient small and nano versions of YOLO.
YOLO v5n is a lightweight version of the YOLO v5 object classification algorithm, Ultralytics (2021). It was the first classification DL classification architecture and the smallest and fastest model in the YOLO v5 family due to striking a balance between model weight and detection accuracy. It has improved detection speed while maintaining a high degree of accuracy compared to previous YOLO versions. Liu et al. (2023).
YOLO v5s achieves a balance between detection accuracy and speed, stood out as another significant choice. The object classification algorithm contains fewer layers and parameters compared to other YOLO v5 models, YOLO v5s prioritised efficiency and ease of training, making it suitable for real-time applications. Despite potential sacrifices in detection accuracy compared to larger models, YOLO v5s is used in various domains, including liquid biopsy of lung cancer and pouring robot object detection Zhao et al. (2023).
Moving to the YOLO v8n, the smallest and fastest model in the YOLO v8 family with 3.2 million parameters, it introduced several architectural improvements over YOLO v5 Selcuk and Serif (2023). YOLO v8n has improved accuracy, faster object detection speed, higher mAP, and the exclusion of anchor boxes during object detection. The enhanced speed and accuracy made YOLO v8n particularly well-suited for real-time detection tasks, aligning seamlessly with the project’s objectives Selcuk and Serif (2023); Wang et al. (2023).
The fourth model, YOLO v8s has more parameters than the other three YOLO models used in this research, 11.2 million parameters—8 million more than YOLO v8n. This increased size should translate to improved performance. The model has more Floating Point Operations Per Second (FLOPS) than its smaller YOLO v8n counterpart Selcuk and Şerif (2023); Selcuk and Serif (2023); Wang et al. (2023).
@@ 3.4 Datasets Pre-processing
The authors have previously generated a public custom underwater crayfish dataset, consisting of 2486 images Machado et al. (2023). This dataset was partitioned into three subsets: 1740 images for the training set, 249 for testing, and 497 images for validation. Additionally, another public custom dataset, also created by the authors, focuses on plastic and comprises 1220 images Machado (2022). Similarly, the plastic dataset was divided into three segments: 854 images for the training set, 122 for testing, and 244 images for validation. A sample is shown in Figure 3.
Figure 3: Sample images from datasets used in training DL models for detecting crayfish and plastic in UK rivers.
The dataset contains images captured from real underwater environments that contain crayfish and a wide variety of plastic debris in different stages of decay and are either clearly visible or covered with varying amounts of overgrowth or other underwater debris. Furthermore, lighting quality and water clarity vary greatly across the dataset which makes the images resemble real-world conditions that the UDEEP prototype is intended to work in.
Image labelling is an important step whereby labels or tags are assigned to images based on their content in order to categorise and organise them in a machine-readable format. The Roboflow computer vision platform, Roboflow (2023) was used to accomplish the labelling task. Using Roboflow, the dataset was labelled with two classes, defined as Crayfish: all signal crayfish or Plastic: all plastic debris/material.
@@ 3.5 Dataset Augmentation
Fine-tuning YOLO models for optimal performance goes beyond merely implementing the algorithm itself. An essential aspect of the labelling process is the utilisation of image augmentation techniques, such as cropping, flipping, rotating, and colour shifting, which can enhance the model’s generalisation and robustness Xu and Matzner (2018). Data augmentation, which is not properly customised, may not yield desired model performance. This section delves into various data augmentation techniques performed on the crayfish and underwater plastic datasets introduced in this article.
  • 1.
    Image Scale augmentation:
    Image Scale augmentation is the action of resizing the input images into various dimensions or scales in order to train the YOLO model on a dataset containing a variety of scales. This adapts the model to objects of differing sizes, which is the case in real-world scenes. The images in the dataset were stretched to 416x416
  • 2.
    Image Flip: Horizontal, Vertical: Images in the dataset were vertically and horizontally flipped. Vertically flipping creates a mirror image where the top becomes the bottom and vice versa. This helps the YOLO model detect objects that may appear upside down in real-world scenarios. Similarly, flip left-right augmentation horizontally flips an image, allowing the model to learn and detect objects from different perspectives. Training on these flipped images enhances the model’s robustness, making it adaptable to accurately detect objects regardless of their orientation
  • 3.
    Cropping: In cases where the subject is obscured by other objects, only a portion of the object can be captured in the photograph. Therefore, we randomly cropped the images, with both the cropping width and height ranging from 0% Minimum Zoom, to 49% Maximum Zoom.
  • 4.
    Mosaic Augmentation: Mosaic augmentation was also applied to the dataset. This is a method that combines multiple images into a single mosaic-like training sample. This technique enhances the YOLO model’s ability to detect objects in complex scenes with overlapping or crowded environments. By training with mosaic-augmented images, the model becomes more adept at handling situations where objects are partially hidden or blend together, ultimately improving its accuracy in challenging scenarios.
  • 5.
    Hue, Saturation, Exposure and Gray-scale: The hue component of images in the dataset was altered in the range between -25° and +25°, saturation altered in the range between -42% and +42%, exposure altered between -22% and +22% and Gray-scale applied to 47% of images.
    Adjusting the hue component mimics diverse lighting conditions, including daylight or artificial lighting. This allows the model to train and recognise objects effectively under a range of illumination settings Applying the Hue, Saturation and exposure augmentation ensures the trained model is robust enough to handle real world lighting exposure, various colour schemes and contrasts
@@ 4 Result Analysis
This section presents the results for training, testing and validation. In the previous section, a classifier was trained with 2300 images in the training set and evaluated its performance on 72 images in the test set. The classifier’s accuracy was assessed based on three measures: training accuracy, test accuracy, and validation accuracy. Training accuracy refers to how accurately the classifier performed on the training set, while test accuracy measures its performance on the test set. Validation accuracy indicates the rate at which the classifier successfully classified newly encountered images in the validation set. Additionally, we considered the recorded losses to evaluate the effectiveness of our approach.
@@
Two functions that return an array of labels was created to produce 18 different confusion matrices. The first function grabs all the true values from the test set. The second function grabs all the predicted labels from each model with the highest mAP value per augmentation. The two arrays of values are then used to plot the confusion matrices. The confusion matrix uses recall equation, precision, and F1-score, equations 2, 3 and 4 respectively.
@@
4.1 Experimental Results on the Crayfish and Plastic Dataset
@@
The figure 4 below present the precision-recall curves scores obtained during the classification task. The graphs all show a high area under the curve, which signifies a combination of high recall and high precision. High precision indicates that the model achieved a low false positive rate, meaning that the classifier accurately identifies crayfish without many false alarms. On the other hand, high recall corresponds to a low false negative rate, implying that the classifier successfully captures the majority of positive results.
When both precision and recall scores are high, it indicates that the classifier is returning accurate and relevant results (high precision) while also capturing a significant portion of all positive results (high recall).
Consequently, an analysis of the results of our training indicate that YOLO v8n achieved the highest precision among the four YOLO architectures achieving a score of 0.92 followed by YOLO v8s that achieved a score of 0.893. YOLO v5s achieved a score of 0.885 and finally YOLO v5n achieving an F1 score of 0.85. These results are shown in the graphs below.
Figure 4: YOLO v5n,YOLO v5s,YOLO v8n and YOLO v8n PR curve.
@@
Figures 5 below presents the F1 confidence score results for the crayfish and plastic classification task. YOLO v5s and YOLO v8s achieved the highest scores at F1 score of 87%. This was followed by YOLO v8n that had an F1 score of 85% and finally YOLO v5s achieving the lowest score of 84%. A summary of these results is also presented in table 1.
Figure 5: YOLO v5n,YOLO v5s,YOLO v8n and YOLO v8n F1-Score.
To measure current usage, Jetson-stats, Bonghi (2019) package was used to analyse the average current used by the board while running the classification models. A python script was adapted to utilise the Jetson-stats library to record the current measurements into a Comma Separated Values (CSV) file stored on the board. The mean inference current consumed during the classification task on the NVIDIA Jetson Nano was calculated based on data collected from classifying 69 images. Measuring power consumption is crucial to guarantee that the device’s power requirements are adequately addressed. Effective power source planning is essential, as it enables the implementation of backup measures for devices operating remotely at the edge. Such proactive measures aim to minimise the potential risk of downtime caused by power outages.
Table 1: Yolo training results.
Class Precision Recall F1 Score Mean Inference Current (mA)
YOLO v5s 0.885 0.92 0.87 3158
YOLO v5n 0.851 0.93 0.84 3277
YOLO v8s 0.917 0.92 0.85 2548
YOLO v8n 0.83 0.92 0.87 2710
Table 1 shows the combined validation results for training both the crayfish and plastics and the mean inference current in mA consumed by the NJN board during inference. The results show that YOLO v8s requires the least amount of current and YOLO v5n requires the most current for inference.
4.2 UDEEP In-Situ underwater Testing
To showcase the effectiveness and benefits of the UDEEP underwater object detection approach employing the four YOLO network models, experiments were conducted in real underwater environments in Rivers within the East of England. The experiment involved fully submerging the UDEEP prototype in the river where crayfish sightings had been reported.
Appropriate permission from the landowners, Health and safety training and licensing had to be obtained. The UDEEP prototype operated fully off-grid due to its electric power source being supplied by rechargeable batteries charged by solar power cells. Figure 6 displays visualisations of the detected objects as part of the experiment results.
Figure 6: UDEEP crayfish and plastic detection results.

### v2

@@ 3 Methodology
The methodology section is divided into four parts. The first part introduces the CED platform workflow, followed by the hardware setup. The third subsection explains the YOLO v5n, YOLO v5s, YOLO v8n and YOLO v8s model architectures and final subsection reviews the dataset augmentation techniques applied to the crayfish and underwater plastic datasets. YOLO was selected for CED object detection task due to its strong performance in terms of speed, accuracy, and deployment efficiency. In contrast to two-stage detectors like Faster R-CNN [ren2015faster], which first generate region proposals and then classify them, YOLO uses a single-stage architecture that directly predicts bounding boxes and class probabilities from the full image in one forward pass [redmon2016you]. These characteristics are particularly important for environmental monitoring applications, where models may need to run on embedded or edge devices with limited computational resources. YOLO ’s lightweight variants (e.g., YOLO v5n, YOLO v8n) can be deployed efficiently on such platforms while still providing reliable detection of small, variably shaped targets like fish and discarded plastic debris  [jocher2020yolov5, glenn2023yolov8]. The source code used in this work is available at https://github.com/denomon/CognitiveEdgeDeviceForRTEMonitoring.
3.1 CED prototype overview
@@
A high level view of the implemented prototype for aquatic monitoring is illustrated in Figure 1. The system features an NJON equipped with an Intel RealSense D435i depth sensor. The NJON is used for accelerating the YOLO v5 and YOLO v8 inference at the edge. The NJON selected is tailored for diverse tasks, including image identification, object detection, data segmentation, and audio processing. The NJON contains a 512-core NVIDIA GPU with 16 Tensor Cores, 6-core Arm Cortex-A78AE v8.2 64-bit Application Processor Unit (APU) and 8GB 128-bit LPDDR5 with 68 GB/s bandwidth.
Figure 1 below shows four distinct CED workflow modules. The first module, Data capture module is responsible for underwater vision data capture using specialised underwater cameras. The videos serve as the primary inputs for the signal crayfish and plastic detection module and are recorded continuously to ensure comprehensive coverage of the surveyed area. The module is also enhanced with sensors for temperature, salinity, pH, dissolved oxygen, and depth to provide vital context for underwater monitoring. Combining sensor data with vision data allows richer, multi-modal analysis of aquatic environments and biodiversity.
The second module, The Signal Crayfish and Plastic Detection & Tracking module handles data extraction, processing, signal crayfish and plastic classification and tracking. Data preparation includes image resizing, scaling, and noise reduction. YOLO pre-trained Deep Learning (DL) models perform inference, identifying and tracking Signal crayfish and plastic objects within the extracted frames. The module summarises sensor data, linking detected objects with corresponding timestamps and environmental parameters. The final stage transmitts the results via network to cloud servers for onward storage and analysis. Sensor metadata, detection results, and positional information are stored for analysis whenever required in the future.
Figure 1: CED monitoring workflow - end-to-end process of underwater monitoring. The system begins by recording underwater videos, followed by data import where camera and sensor metadata are stored and video frames extracted. The detection and tracking module then performs real-time identification of Signal Crayfish and plastic waste across frames. Finally, the data archiver formats metadata such as timestamps, geo-location, and depth, storing both detection and sensor data for later export and analysis, enabling biodiversity monitoring and pollution assessment.
@@ 3.2 Network Architecture
Four object classification network architectures were used from the YOLO family used in this research project, driven by the necessity to perform object detection at the edge where processing and power resources are constrained. The choice leaned towards the compact and efficient small and nano versions of YOLO. YOLO v5n is a lightweight version of the YOLO v5 object classification algorithm, [ultralytics]. It was the first classification DL classification architecture and the smallest and fastest model in the YOLO v5 family due to striking a balance between model weight and detection accuracy. It has improved detection speed while maintaining a high degree of accuracy compared to previous YOLO versions [liu-2023].
YOLO v5s achieves a balance between detection accuracy and speed, stood out as another significant choice. The object classification algorithm contains fewer layers and parameters compared to other YOLO v5 models, YOLO v5s prioritised efficiency and ease of training, making it suitable for real-time applications. Despite potential sacrifices in detection accuracy compared to larger models, YOLO v5s is used in various domains, including liquid biopsy of lung cancer and pouring robot object detection [zhao-2023].
Moving to the YOLO v8n, the smallest and fastest model in the YOLO v8 family with 3.2 million parameters, it introduced several architectural improvements over YOLO v5 [selcuk2023comparison]. YOLO v8n has improved accuracy, faster object detection speed, higher mAP, and the exclusion of anchor boxes during object detection. The enhanced speed and accuracy made YOLO v8n particularly well-suited for real-time detection tasks, aligning seamlessly with the project’s objectives [selcuk2023comparison, wang2023uav].
The fourth model, YOLO v8s has more parameters than the other three YOLO models used in this article, 11.2 million parameters, 8 million more than YOLO v8n. The model also has more Floating Point Operations Per Second (FLOPS) than its smaller YOLO v8n counterpart [selcuk-2023, selcuk2023comparison, wang2023uav] resulting in improved performance than the three YOLO models used in this article.
@@ 3.3 Datasets Pre-processing
The authors have previously generated a public custom underwater crayfish dataset, consisting of 2486 images [machado_2023]. The proposed dataset was partitioned into three subsets: 1740 images for the training set, 249 for testing, and 497 images for validation. Additionally, another public custom dataset, also created by the authors, focuses on plastic and comprises 1220 images [machado_2022]. Similarly, the plastic dataset was divided into three segments: 854 images for the training set, 122 for testing, and 244 images for validation. A sample is shown in Figure 2.
The dataset contains images captured from real underwater environments that contain crayfish and a wide variety of plastic debris in different stages of decay and are either clearly visible or covered with varying amounts of overgrowth or other underwater debris. Furthermore, lighting quality and water clarity vary greatly across the dataset which makes the images resemble real-world conditions that the CED prototype is intended to work in.
Image labelling is an important step whereby labels or tags are assigned to images based on their content in order to categorise and organise them in a machine-readable format. The Roboflow computer vision platform, [roboflow] was used to accomplish the labelling task. Using Roboflow, the dataset was labelled with two classes, defined as Crayfish: all signal crayfish or Plastic: all plastic debris/material.
Figure 2: A compilation showcasing a variety of underwater images that are included in the underwater plastic and signal crayfish dataset. A) A clear view of a sandy or gravelly bottom. There’s a piece of decaying plastic material. B) A large, brownish plastic material in deeper or murkier water, obscuring the view causing limited visibility of the surrounding environment. C) A partially submerged plastic debris D) A close-up shot of a signal crayfish being held by a hand with claws and legs clearly visible. E) A view of a sandy or gravelly bottom with a juvenile signal crayfish partially visible. F) Numerous juvenile signal crayfish swimming amongst rocks and shells on the bottom.
@@ 3.4 Dataset Augmentation
Fine-tuning YOLO models for optimal performance goes beyond merely implementing the algorithm itself. An essential aspect of the labelling process is the utilisation of image augmentation techniques, such as cropping, flipping, rotating, and colour shifting, which can enhance the model’s generalisation and robustness [xu-2018]. Data augmentation, which if not properly customised, may not yield desired model performance. The various data augmentation techniques performed on the crayfish and underwater plastic datasets introduced in this article are discussed below.
Image Scale augmentation is the action of resizing the input images into various dimensions or scales in order to train the YOLO model on a dataset containing a variety of scales. This adapts the model to objects of differing sizes, which is the case in real-world scenes. The images in the dataset were stretched to 416x416 to ensure consistency during training and inference.
Images in the dataset were vertically and horizontally flipped. Vertically flipping creates a mirror image where the top becomes the bottom and vice versa. Vertically flipping helps the YOLO model detect objects that may appear upside down in real-world scenarios. Similarly, flip left-right augmentation horizontally flips an image, allowing the model to learn and detect objects from different perspectives. Training on these flipped images enhances the model’s robustness, making it adaptable to accurately detect objects regardless of their orientation
In cases where the subject is obscured by other objects, only a portion of the object can be captured in the photograph. Therefore, we randomly cropped the images, with both the cropping width and height ranging from 0% Minimum Zoom, to 49% Maximum Zoom.
Mosaic augmentation, a method that combines multiple images into a single mosaic-like training sample was also applied to the dataset. Mosaic augmentation technique enhances the YOLO model’s ability to detect objects in complex scenes with overlapping or crowded environments. By training with mosaic-augmented images, the model becomes more adept at handling situations where objects are partially hidden or blend together, ultimately improving its accuracy in challenging scenarios.
The hue component of images in the dataset was altered in the range between -25° and +25°, saturation altered in the range between -42% and +42%, exposure altered between -22% and +22% and Gray-scale applied to 47% of images. Adjusting the hue component mimics diverse lighting conditions, including daylight or artificial lighting. Altering the hue component allows the model to train and recognise objects effectively under a range of illumination settings Applying the Hue, Saturation and exposure augmentation ensures the trained model is robust enough to handle real world lighting exposure, various colour schemes and contrasts
@@ 4 Result Analysis
The results for training, testing and validation are discussed in this section. In the previous section, a classifier was trained with 2594 images in the training set and evaluated its performance on 371 images in the test set. The classifier’s accuracy was assessed based on three measures: training accuracy, test accuracy, and validation accuracy. Training accuracy refers to how accurately the classifier performed on the training set, while test accuracy measures its performance on the test set. Validation accuracy indicates the rate at which the classifier successfully classified newly encountered images in the validation set. Additionally, we considered the recorded losses to evaluate the effectiveness of our approach.
@@
Two functions that return an array of labels were created to produce 18 different confusion matrices. The first function grabs all the true values from the test set. The second function grabs all the predicted labels from each model with the highest mAP value per augmentation. The two arrays of values are then used to plot the confusion matrices. The confusion matrix, which summarises the counts of True Positives (TP), True Negatives (TN), False Positives (FP), and False Negatives (FN), provides the necessary values for calculating recall (2), precision (3), and F1-score (4).
@@
4.1 Training Results
@@
The figure 3 below presents the precision-recall curve scores obtained during model training. The graphs all show a high area under the curve, which signifies a combination of high recall and high precision. High precision indicates that the model achieved a low false positive rate, which means that the classifier accurately identified crayfish without many false positives. however, high recall corresponds to a low false negative rate, implying that the classifier successfully captures the majority of positive results. When both precision and recall scores are high, it indicates that the classifier is returning accurate and relevant results (high precision) while also capturing a significant portion of all positive results (high recall). Consequently, an analysis of the results of our training indicate that YOLO v5s achieved the highest precision among the four YOLO architectures achieving a score of 0.93 followed by YOLO v8s that achieved a score of 0.91. YOLO v8n achieved a score of 0.90 and finally YOLO v5n achieving an precision score of 0.88. These results are shown in table 1 and figures 3 and 4 below.
Figure 3: PR curves for YOLO v5n, YOLO v5s, YOLO v8n and YOLO v8s models on crayfish and plastic detection. The light blue and orange curves represent the crayfish and plastic classes while the bold blue curve indicates the overall mAP at IoU 0.5. YOLO v5s achieves the highest mAP (0.900), followed closely by YOLO v8s (0.887).
@@
Figures 4 below presents the F1 confidence scores obtained during the training exercise. YOLO v5s and YOLO v8s achieved the highest scores at F1 score of 0.87. This was followed by YOLO v8n that had an F1 score of 0.85 and finally YOLO v5s achieving the lowest score of 0.84. A summary of these results is also presented in table 1.
Figure 4: F1-score vs. confidence threshold curves for YOLO v5n, YOLO v5s, YOLO v8n, YOLO v8s models. The light blue and orange curves represent the crayfish and plastic classes while the bold blue curve indicates the overall F1-score. YOLO v8s achieves the highest F1-score of 0.87 at a confidence threshold of 0.626, followed closely by YOLO v8n 0.85 at a confidence threshold of 0.541.
Table 1: Comparison of final validation performance across four YOLO object‐detection variants on our crayfish and aquatic plastic dataset. Bold indicates the best performance in each column.
Model Precision Recall F1 Score mAP@0.5
YOLO v5s 0.93 0.83 0.87 0.90
YOLO v5n 0.88 0.81 0.84 0.87
YOLO v8s 0.91 0.84 0.87 0.88
YOLO v8n 0.90 0.80 0.85 0.86
4.2 Convergence analysis
We evaluated the convergence stability of YOLO v5s, YOLO v5n, YOLO v8s, and YOLO v8n based on mAP@0.5 and validation box loss over 580 training epochs. YOLO v5s achieved the highest final mAP@0.5 (0.900), followed closely by YOLO v8s (0.887). YOLO v5n and YOLO v8n showed comparable but slightly lower final mAP values (0.874 and 0.873, respectively). The small models (YOLO v5s and YOLO v8s) converged more quickly and smoothly, with less fluctuation in performance across epochs.
The loss trends of the validation boxes aligned with the mAP curves, YOLO v8s recorded the lowest and most stable validation loss, indicating efficient localisation, while YOLO v5n exhibited higher and more erratic loss, consistent with its slower and less stable convergence. Overall, YOLO v5s demonstrated the best balance between high accuracy and stable training, making it the most effective model for the CED detection task, with YOLO v8s also performing robustly. The results are shown in figure 5 below.
Figure 5: Convergence analysis of YOLO v5s, YOLO v5n, YOLO v8s, and YOLO v8n over 580 training epochs. The top plot shows mAP@0.5 progression, where YOLO v5s converges fastest and achieves the highest detection accuracy. The bottom plot shows validation box loss, with YOLO v8s exhibiting the lowest and most stable loss, indicating superior localization performance. Overall, YOLO v8s demonstrates the best convergence speed and training stability among the evaluated models.
5 Model performance and power consumption on the NJON
Measuring power consumption and inference time is crucial to ensure that device inference speed and power consumption requirements are adequately addressed. Effective power source planning is essential as it enables the implementation of power backup measures for devices operating remotely at the edge. Such proactive measures aim to minimise the potential risk of downtime caused by power outages. To measure power consumption on the the NJON, Jetson-stats, [jetsonstats] python library was used to analyse the average power consumed by the NJON platform while performing inference using the four YOLO classification models, one at a time. CPU, GPU Inference time and power consumption measurements were the key metrics recorded and the results shown in table 2 are discussed below. Energy consumption was estimated by multiplying the inference time by the change in total board power (VDD_IN), recorded immediately before and after inference using Jetson-stats. This approach provides a practical approximation for comparing model energy efficiency on the Jetson platform because we are able to capture the change in power consumption during the inference period. We report energy consumption rather than power consumption, as model inference times vary significantly. This provides a fairer comparison of overall resource efficiency on the NJON. The NJON platform was selected to evaluate its potential for real-time environmental monitoring, especially in aquatic ecosystems where accurate classification of invasive species is critical to the conservation of biodiversity. The results are discussed in sections 5.1 and 5.2 below.
Table 2: Inference time and power consumption for four YOLO models on CPU and GPU. YOLO v8n highlighted in green colour below achieved the best inference time (fastest) and consumed the least amount of energy per inference on the GPU. GPU consistently achieves best inference time and required lower energy per inference compared to CPU.
Model Device Inference Time(s) Power(W) Energy(J)
YOLO v8n GPU 0.2914 0.32 0.093248
YOLO v8s GPU 0.3748 1.19 0.446012
YOLO v5s GPU 0.4342 0.44 0.191048
YOLO v5n GPU 0.4829 0.83 0.400807
YOLO v8n CPU 0.5947 1.24 0.737428
YOLO v5n CPU 0.8579 2.43 2.084697
YOLO v5s CPU 1.3083 1.03 1.347549
YOLO v8s CPU 1.3125 0.87 1.141875
5.1 Inference Time
The four YOLO-based models showed significant differences in their inference times depending on the device (CPU vs. GPU) used as shown in figure 6 below. When executed on the GPU, YOLO v8n demonstrated the fastest inference time, completing classification in 0.29 seconds. Conversely, YOLO v5n had the longest inference time on the GPU, requiring 0.48 seconds. In contrast, CPU executions were notably slower across all four models. The YOLO v8n model took 0.59 seconds on the CPU, while YOLO v8s was the slowest with a CPU inference time of 1.31 seconds. These results highlight the efficiency of the GPU in handling real-time object detection tasks, particularly for embedded systems like the NJON. Based on these findings, YOLO v8n proved to be the fastest model for real-time inference on the NJON platform.
Figure 6: Bar chart showing the inference time of four YOLO models on GPU and CPU, ordered by the lowest inference time. The results show that GPU achieves faster inference time compared to the CPU with YOLO v8n achieving the lowest results on the GPU and YOLO v8s taking longest inference time CPU
5.2 Energy Consumption Measurements
To assess the energy efficiency of the deployed YOLO models on the NVIDIA Jetson platform, we estimated the energy consumed during inference using power readings obtained via Jetson-stats. The total NJON board power consumption was monitored using the VDD_IN values reported before and after inference. The increase in power consumption during inference was computed as:
$\Delta\text{VDD}_{\text{IN}}=\text{VDD}_{\text{IN, After}}-\text{VDD}_{\text{IN, Before}}$ (5)
This method provides a consistent way to compare the relative energy costs of different models, under the assumption that the primary power increase during the short inference window is attributable to the model execution.The inference energy consumption $E$ (in J) was then estimated using:
$E=\Delta\text{VDD}_{\text{IN}}\times\text{Inference Time (s)}$ (6)
From the results shown in figure 7 below, it was observed that the GPU generally required less energy than the CPU across all four custom trained YOLO models. The YOLO v8n model, which had the fastest inference time of 0.29 seconds on the GPU, also demonstrated the lowest GPU energy consumption at 0.09J and YOLO v8s had the highest GPU energy consumption at 0.45J per inference. On the CPU, energy consumption ranged between 0.74J and 2.08J, with YOLO v8n consuming the least CPU power of 0.74J and YOLO v5n consuming the highest amount of power at 2.08J. These findings indicate that YOLO v8n is best suited for power-limited edge environments, as it consumes the least amount of power among the four YOLO variants.
The results also indicate that GPU execution provides significant advantages in terms of both inference time and power consumption for deep learning tasks like signal crayfish and plastic classification. The YOLO v8n model, which exhibited the best balance of performance and low power consumption, is the most suitable option for real-time deployment on the NJON platform. On the other hand, CPU execution, while offering flexibility in deployment scenarios, generally results in higher power consumption and longer inference time. These findings are crucial for optimising the performance of edge devices in resource-constrained environments, where both power and speed are essential for practical applications.
Figure 7: NJON GPU and CPU power consumption measurements while performing a classification task using four pre-trained YOLO models. GPU power consumption is significantly lower than CPU power consumption with YOLO v8n consuming the least amount at 0.09J on the GPU and YOLO v5n consuming the most at 2.08J on the CPU
5.3 Stability
To evaluate the performance of different YOLO models, we conducted a series of tests to evaluate the inference time, power consumption, and robustness. The models tested are YOLO v5n, YOLO v5s, YOLO v8n, and YOLO v8s, with a focus on three key areas: stability, robustness to Gaussian noise and robustness to motion blur which are common challenges facing underwater environments. The tests were executed on the GPU due to prior tests that confirmed GPU execution provides significant advantages in terms of both inference time and power consumption for deep learning tasks.
Stability testing involved evaluating the consistency of inference time, power consumption, and confidence stability when the same input image was used repeatedly. The stability of the models was measured by calculating the mean and standard deviation of inference times and power consumption over several runs. YOLO v5n exhibited a mean inference time of 0.375 seconds with a relatively high standard deviation (0.495), indicating some variability. YOLO v5s showed a higher mean inference time of 0.577 seconds, with even more significant variability (0.684). YOLO v8n was the fastest, with a mean inference time of 0.250 seconds and a moderate standard deviation of 0.532. YOLO v8s had a mean inference time of 0.459 seconds, with a standard deviation of 0.695, slightly more variability than YOLO v8n.
5.4 Robustness to Gaussian noise and motion blur
For robustness, the performance of each of the four YOLO models was tested under two conditions: Gaussian noise and motion blur. Gaussian noise is commonly used to simulate sensor noise and environmental distortions in underwater imaging, which can arise due to low-light conditions or electronic interference [zhangWang2021]. Motion blur mimics the effects of fish movement and camera instability, helping assess the robustness of the YOLO models in dynamic aquatic environments [chen2020]. The robustness test with Gaussian noise simulates poor lighting or sensor artifacts in real-world scenarios. YOLO v5n showed a slight increase in inference time (0.394 seconds), with a high standard deviation (0.671), but power consumption remained minimal. YOLO v5s exhibited a significant drop in mean inference time (0.251 seconds), but its standard deviation increased considerably (0.501), indicating some instability. YOLO v8n performed well, with the fastest inference time (0.228 seconds) and a low deviation (0.496), showcasing strong robustness. YOLO v8s, on the other hand, was the slowest at 0.594 seconds, with a very high standard deviation (1.065), suggesting instability under noisy conditions.
In the motion blur test, which simulates detection in dynamic settings, YOLO v5n showed a decrease in inference time (0.206 seconds) but had a high deviation (0.470). YOLO v5s performed inconsistently, with a high inference time (0.600 seconds) and significant variability (1.366). YOLO v8n performed exceptionally well, with the lowest mean inference time of 0.056 seconds and minimal variation (0.045), making it highly robust to motion blur. YOLO v8s had a high mean inference time of 0.772 seconds, along with a high standard deviation of 0.859, indicating less robustness in this scenario. The table below summarises the results of stability and robustness to noise (Gaussian and motion blur), highlighting the mean and standard deviation of inference times and power consumption for each model and test scenario.
Model Test Type Mean Inference Time (s) Mean GPU +CPU Power (W) Mean System On Chip (SOC) Power (W) Mean inference Confidence YOLO v5n Stability 0.2210 0.0676 0.0314 0.7993 YOLO v5n Robustness (Gaussian Noise) 0.0687 0.0160 0.0160 0.6803 YOLO v5n Robustness (Motion Blur) 0.0719 0.0161 0.0201 0.5265 YOLO v5s Stability 0.1310 0.0325 0.0202 0.8728 YOLO v5s Robustness (Gaussian Noise) 0.0894 0.0285 0.0081 0.8579 YOLO v5s Robustness (Motion Blur) 0.0912 0.0161 0.0201 0.8454 YOLO v8n Stability 0.0928 0.0119 0.0160 0.8448 YOLO v8n Robustness (Gaussian Noise) 0.0720 0.0081 0.0001 0.7757 YOLO v8n Robustness (Motion Blur) 0.0767 0.0160 0.0081 0.7024 YOLO v8s Stability 0.1226 0.0647 0.0200 0.8542 YOLO v8s Robustness (Gaussian Noise) 0.0905 0.0325 0.0080 0.7128 YOLO v8s Robustness (Motion Blur) 0.0944 0.0244 0.0121 0.8041
Table 3: Performance evaluation of different YOLO models under various test conditions ( mean inference time and power consumption (watts), and mean confidence). Best values are highlighted in bold, worst in italics.